Christine and Jessica leave from the same spot at the same time and begin jogging in opposite directions. At the end of an 1 1/2 hours they are 15 miles apart. Find Christine's average speed if her average speed is 2 miles per hour more than Jessica's average speed.
Answer by CharlesG2(834)

Christine and Jessica leave from the same spot at the same time and begin jogging in opposite directions. At the end of an 1 1/2 hours they are 15 miles apart. Find Christine's average speed if her average speed is 2 miles per hour more than Jessica's average speed.

distance D = rate R * time T
let R = Jessica's average speed
let R + 2 = Christine's average speed
let T = 1 1/2 = 1.5 hours
let D = total distance of 15 miles
15 = R * 1.5 + (R + 2)1.5
15 = 1.5R + 1.5R + 3
12 = 3R
4 = R
Jessica's average speed is 4 mph
Christine's average speed is 6 mph
Jessica's distance = 4 * 1.5 = 6 miles
Christine's distance = 6 * 1.5 = 9 miles
6 miles + 9 miles = 15 miles